# STC-ViT: Spatio Temporal Continuous Vision Transformer for Weather Forecasting

### v2

@@ Abstract
Operational weather forecasting system relies on computationally expensive physics-based models. Recently, transformer based models have shown remarkable potential in weather forecasting achieving state-of-the-art results. However, transformers are discrete models which limit their ability to learn the continuous spatio-temporal features of the dynamical weather system. We address this issue with STC-ViT, a Spatio-Temporal Continuous Vision Transformer for weather forecasting. STC-ViT incorporates the continuous time Neural ODE layers with multi-head attention mechanism to learn the continuous weather evolution over time. The attention mechanism is encoded as a differentiable function in the transformer architecture to model the complex weather dynamics. We evaluate STC-ViT against a operational Numerical Weather Prediction (NWP) model and several deep learning based weather forecasting models. STC-ViT performs competitively with current data-driven methods in global forecasting while only being trained at lower resolution data and with less compute power.
@@ 1 Introduction
Operational weather forecasting systems rely on physics based Numerical Weather Prediction (NWP) models. While highly accurate, these system tend to be extremely slow with aggregated approximation errors Palmer et al. (2005) and requiring high computational resources usually hundreds of thousands of GPU nodes which is unfeasible Bauer et al. (2020) Andersson (2022). The recent transformers based Deep Learning (DL) architectures Nguyen et al. (2023a); Kurth et al. (2023); Nguyen et al. (2023b), have achieved impressive forecasting results outperforming the NWP models.
Figure 1: Conceptual representation of continuous dynamics of weather system entailing that the rate of change of air density in a given volume is balanced by the flux of air mass entering or leaving the volume.
Weather forecasting is based on the Continuity Equation $\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\textbf{u})=0$, which ensures the mass of air in the atmosphere remains constant over time, where $\rho$ is the air density, t is time, u is the velocity vector of the air, which has components u, v, and w representing the wind speeds in the x (east-west), y (north-south), and z (vertical) directions, respectively and $\nabla\cdot(\rho\textbf{u})$ represents the divergence of the mass flux.
However, transformers only process discrete sequences of input data and are fundamentally discrete in nature Fonseca et al. (2023) which limit their ability in modelling the continuous evolving dynamics of weather system.
@@
In this work, we address the non-continuity problem in Vision Transformer (ViT) which arise in weather forecasting systems where data is present in discrete form. The distinctive and continuously evolving characteristics of weather data pose significant challenge in producing accurate forecasts. The correlation within weather information observed at different time steps is continuously evolving due to the continuous nature of data and processing it as discrete information greatly impacts the forecasting accuracy.
To tackle the temporal dynamics of a system, Neural Ordinary Differential Equation (NODE) Chen et al. (2018) was introduced. NODE is based on the concept of continuous depth modelling to learn smooth solutions from irregular time series data. However, while capturing the complex dependencies between observations, the iterative approach of Neural ODE’s can lead to escalating errors as the number of iterations increases Chen et al. (2018). Moreover, despite being computationally efficient and accurate, Neural ODEs are only limited to temporal data and fails to learn the continuous features of spatio-temporal dynamical systems such as weather forecasting.
@@
To address these problems, we build upon the core ideas of Neural ODE Chen et al. (2018) and ViT Dosovitskiy et al. (2020). We propose STC-ViT which leverages the continuous learning paradigm to effectively learn the complex spatio-temporal changes even from weather data recorded at coarser resolution. The idea is to parameterize the attention mechanism by converting it into a differentiable function. Continuous temporal attention is calculated sample-wise and combined with the patch wise spatial attention to learn the spatio-temporal mapping of weather variables in the embedding space of the vision transformer. Furthermore, we add derivation as a pre-processing step to prepare the discrete data for continuous model and explore the role of normalization in continuous modelling.
Normalization is a well established concept in the field of deep learning especially in vision transformers where pre and post layer normalization optimize gradient propagation under different settings. The influence of different normalization techniques in neural ODEs was also explored by Gusak et al. however, the role of normalized derivatives is still poorly understood in complex continuous spatio-temporal data dependant forecasting systems. We empirically investigate the influence of normalization in forecasting different weather variables. We observed that normalizing the continuous evolution of hidden layers yields different results for each predicted weather variable. Empirical results show the ability of the proposed STC-Vit to model complex continuous spatio-temporal dynamic systems effectively and achieve state-of-the-art performance in global weather forecasting.
@@
In this paper we focus on the following research question: Can we design a data-driven weather forecasting system that is computationally efficient, learns the continuous latent representation and achieves precision levels commensurate with Numerical Weather Prediction (NWP) model. In summary, our contributions are as follows:
@@
Spatio-temporal continuous attention: We propose continuous attention to effectively learn and interpolate the spatio-temporal information for weather forecasting.
@@
Data pre-processing for continuous systems: We introduce derivation as a pre-processing step to ensure better feature extraction for continuous spatio-temporal models.
Normalization in continuous modelling: We explore the role of normalization in STC-ViT and conduct extensive experiments to analyse our methodology.
@@ 2 Background
2.1 Attention in Vision Transformers.
The attention mechanism plays a crucial role in capturing long-term spatial dependencies. While the concept of attention in vision has been discussed in literature for decades, it gained popularity with the introduction of Vision transformers Dosovitskiy et al. (2020). The attention mechanism enables the model to focus on different parts of the input image simultaneously, capturing various features and their interactions at different positions. It allows the model to weigh the importance of different input features dynamically. In a transformer, attention is calculated using a set of queries $(Q)$, keys $(K)$, and values $(V)$, typically derived from the input features. The attention mechanism can be described mathematically as:
$Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (1)
where Q, K, and V are matrices formed from all queries, keys, and values, respectively, $d_{k}$is the dimensionality of the keys and queries, and the division by $\sqrt{d_{k}}$ is a scaling factor to deal with gradient vanishing problems during training.
In practice, multi-head attention is used in which the attention mechanism is run several times in parallel allowing the model to jointly attend to information from different representation subspaces at different positions enhancing the model’s ability to represent complex dependencies.
@@ 2.2 Neural ODE
Neural ODEs are the continuous time models which learn the evolution of a system over time using Ordinary Differential Equations (ODE) Chen et al. (2018). The key idea behind Neural ODE is to model the derivative of the hidden state using a neural network. Consider a hidden state $h(t)$ at time $t$ In a traditional neural network, the transformation from one layer to the next could be considered as moving from time $t$ to $t+1$. In Neural ODEs, instead of discrete steps, the change in $h(t)$ over time is defined by an ordinary differential equation parameterized by a neural network:
@@
From ResNets to Neural ODEs To see the connection between ResNets and Neural ODEs, consider a ResNet with layers updating the hidden state as:
@@ 3 Related Work
WeatherBench Rasp et al. (2020) provides a benchmark platform to evaluate data-driven systems for effective development of weather forecasting models. Current state-of-the-art DL based models are mostly based on Graph Neural Networks (GNNs) and Transformer based models. Keisler Keisler (2022) implemented a message passing GNN based forecasting model which was further extended by GraphCast Lam et al. (2023) which implemented multi-mesh GNN to achieve state-of-the-art results. Pangu-Weather Bi et al. (2023) a 3D transformer based model was introduced which was designed to encode 3D earth pressure levels to reduce the forecasting errors. FourCastNet (FCN) Kurth et al. (2023) used a combination of Fourier Neural Operator (FNO) and ViT and was reported to be 80,000 times faster than NWP models. ClimaX Nguyen et al. (2023a) was introduced as a foundation model for weather and climate forecasting. The authors used Vision Transformer (ViT) as a backbone for spatio-temporal forecasting on number of tasks i.e. global forecasting, regional forecasting and seasonal forecasting each with a lead time of several hours, days and months.
Another data-driven model FengWu Chen et al. (2023a) showed remarkable performance on medium range weather forecasting. It uniquely approaches forecasting through a multi-modal and multi-task perspective, utilizing a deep learning architecture with model-specific encoder-decoders and a cross-modal fusion Transformer. Followed by FengWu, FuXi Chen et al. (2023b) emerged as high resolution forecasting system operating at $0.25^{\circ}$. FuXi integrates the data cube embedding with U-Transformer to produce 15-day forecasts. Recently, Stormer Nguyen et al. (2023b) based on transformer model emerged as the state-of-the-art weather forecasting model outperforming PanguWeather and GraphCast. While being highly accurate and showcasing remarkable scaling capabilities, these models are discrete space-time models and do not account for the continuous dynamics of weather system except for ClimODE Verma et al. (2023). ClimODE uses Neural ODE to incorporate a continuous-time process that models weather evolution and advection, enabling it to capture the dynamics of weather transport across the globe effectively, however it currently yields less precise forecast results compared to state-of-the-art models, offering significant potential for further enhancements.
@@ 4 Methodology
Problem Formulation. Consider a model $f$ receives weather data as input of the form $X^{N\times W\times H}$ at time $t$ and predicts the weather information at time $t+\triangle t$ as shown in equation 5 where $N$ is the number of weather variables such as temperature, wind speed, pressure, etc., $H\times W$ refers to the spatial resolution of the variable.
@@
where $V(x,y,t)$ is pixel value of weather variable at time t and $V(x,y,t-1)$ is the pixel value of same variable at t-1. Figure 2 shows the graphical representation of the pre-processing step.
Figure 2: Temporal derivatives of an image time series capture the changes in pixel values over time. The process involves calculating the rate at which the intensity values of each pixel change from one time frame to the next.
Normalized Derivatives. Normalization of derivatives is one of the crucial steps to encode stability in the behavior of deep learning model. While the effects and results of different normalization techniques in neural ODEs has been explored before Gusak et al. (2020), it only looks at normalization as layers in the neural network architecture. We extend the idea by investigating the role of normalization when directly applied to derivatives. Normalizing derivatives involves scaling the derivative values so that they fall within a specific range or conform to a standard distribution. This normalization process can help improve the stability and performance of subsequent analysis. We applied two normalization most common techniques namely min-max and z-score normalization followed by the pre-derivation layer to showcase their benefits in continuous systems.
@@ 4.1 Spatio-Temporal Continuous Vision Transformer
Attention is one of the key components in the vision transformer architecture. It is based on the idea of determining the most crucial parts of image in the final prediction stage. Despite being successful in variety of tasks, ViT is still restrictive in its ability to learn informative embedding for a highly dynamical system like weather forecasting. We design multiple attention paths Temporal Continuous Attention (TCA) and Spatial Attention (SA) to model the the continuous evolving state of weather variables. The detailed architecture of STC-ViT is shown in figure 3.
Temporal Continuous Attention (TCA). First, we employ sample-wise attention to calculate the contextual embedding space for each weather variable evolving with time. This step ensures that in addition to spatial attention, the model also attends to the temporal evolution of the same variable across different samples within the batch. This temporal transformation is learned by assigning each variable with its own Query, Key, and Value across each input sample, similar to how it’s done within a single image. The attention mechanism calculates the attention scores between the variables of different samples (across the same variable positions).
This modification allows the model to capture relationships or dependencies between the same weather variables across different instances. It proved to be beneficial in weather forecasting scenario where the model is able to map the continuous evolving features of each weather variable. To ensure continuous learning, we introduce derivatives in the sample wise attention mechanism. Differential equations represent the dynamics of the physical system with respect to time and cater to the missing values of data. We combined the attention mechanism with the differential equation learning paradigm to model the change of the atmosphere for both the spatial and temporal features. Moreover, this approach removes the limitation of modelling complex physics equations in NWP models. Furthermore, instead of having predictions at particular time stamp only, STC-ViT learns the transitional changes from one time step to another which is important for capturing unprecedented changes in weather.
@@
To compute TCA, we calculate derivatives of similarity across same variables of each data sample. Assume we have two input samples of size $(N\times W\times H)$ denoted as $X_{0}$ at time $t_{0}$ and $X_{1}$ at time $t_{1}$ where $N$ is the number of weather variables such as temperature, wind speed, pressure, etc., $H\times W$ refers to the spatial resolution of the variable. Each variable has its own Q, K, and V tensors across both samples. Given $Q_{t_{0},i}$ and $K_{t_{1},j}$ are the Query and Key representations for the $i_{th}$ variable at time $t_{0}$ and $j_{th}$ variable at time $t_{1}$. Continuous attention is computed as follows:
$\frac{d}{dt}(Q_{t_{0},i}).(K_{t_{1},j})=Q_{t_{0},i}\frac{d(K_{t_{1},j})}{dt}+K_{t_{1},j}\frac{d(Q_{t_{0},i})}{dt}$ (7)
$Q_{t_{0},i}\frac{d(K_{t_{1},j})}{dt}+K_{t_{1},j}\frac{d(Q_{t_{0},i})}{dt}=Y$ (8)
$Att(Q_{t_{0},i},K_{t_{1},j})=Softmax(\frac{Y}{\sqrt{d_{k}}})$ (9)
$TCA_{t_{0},t_{1}}=\sum\limits^{N}_{k=1}Att(Q_{t_{0},i},K_{t_{1},j})\cdot V_{t_{1},j}$ (10)
@@
The resulting output represents the attention weighted sum of values for similar variables across input samples at time $t_{0}$ and $t_{1}$. This information captures relationships or dependencies between corresponding variables across the input samples at different time steps. This process calculates attention between similar variables across input samples at all time steps, allowing the model to capture relationships or interactions between variables across the entire sequence of input samples.
Spatial Attention (SA). For spatial attention, we simple calculate traditional attention scores along patches of the same image and obtain output as shown in equation 1 and call it SA.
@@
$STC=concat(TCA,SA)$ (11)
@@
Figure 3: Whole prediction pipeline of STC-ViT as the model receives spatio-temporal weather information and passed through pre-derivation and normalization steps and input to transformer encoder where TCA and SA learns the continuous weather features and outputs the prediction

### v3

@@ Abstract
Operational weather forecasting system relies on computationally expensive physics-based models. Recently, transformer based models have shown remarkable potential in weather forecasting achieving state-of-the-art results. However, transformers are discrete and physics-agnostic models which limit their ability to learn the continuous spatio-temporal features of the dynamical weather system. We address this issue with STC-ViT, a Spatio-Temporal Continuous Vision Transformer for weather forecasting. STC-ViT incorporates the continuous time Neural ODE layers with multi-head attention mechanism to learn the continuous weather evolution over time. The attention mechanism is encoded as a differentiable function in the transformer architecture to model the complex weather dynamics. Further, we define a customised physics informed loss for STC-ViT which penalize the model’s predictions for deviating away from physical laws. We evaluate STC-ViT against operational Numerical Weather Prediction (NWP) model and several deep learning based weather forecasting models. STC-ViT, trained on 1.5-degree 6-hourly data, demonstrates computational efficiency and competitive performance compared to state-of-the-art data-driven models trained on higher-resolution data for global forecasting.
@@ 1 Introduction
Operational weather forecasting systems rely on physics based Numerical Weather Prediction (NWP) models. While highly accurate, these system tend to be extremely slow with aggregated approximation errors and requiring high computational resources (Palmer et al., 2005; Andersson, 2022).
Weather forecasting can be formulated as a continuously evolving physical phenomenon which can be typically expressed as time-dependent partial differential equations (PDEs). These PDEs capture the dynamic evolution of atmospheric variables over time, enabling the simulation of complex weather patterns and phenomena. Encoding these physical priors leads to more interpretable and physically consistent forecasting models (Couairon et al., 2024). Recent works have shown that using deep learning models as surrogates to model complex multi-scale spatio-temporal phenomena can lead to training-efficient models (Gupta & Brandstetter, 2022). Vision Transformer by Dosovitskiy et al. (2020) has emerged one of such models (Lessig et al., 2023; Nguyen et al., 2023b).
@@
However, transformers only process discrete sequences of input data and are fundamentally discrete in nature Fonseca et al. (2023) which limit their ability in modelling the continuous evolving dynamics of weather system. Additionally, these models ignore the fundamental physical laws of our atmosphere which is essential for faithful weather modelling. In this work, we address the non-continuity problem in Vision Transformer (ViT) which arise in weather forecasting systems where data is present in discrete form. The distinctive and continuously evolving characteristics of weather data pose significant challenge in producing accurate forecasts. Additionally, we penalise the model for not following physical dynamics in the form a customized physics informed loss function.
@@
We build upon the core ideas of Neural Ordinary Differential Equation (NODE) by Chen et al. (2018) and ViT (Dosovitskiy et al., 2020). We propose STC-ViT which leverages the continuous learning paradigm to effectively learn the complex spatio-temporal changes even from weather data recorded at coarser resolution. The idea is to parameterize the attention mechanism by converting it into a differentiable function. Continuous temporal attention is calculated sample-wise and combined with the patch wise spatial attention to learn the spatio-temporal mapping of weather variables in the embedding space of the vision transformer. Furthermore, we add derivation as a pre-processing step to prepare the discrete data for continuous model and explore the role of normalization in continuous modelling.
@@
In this paper we focus on the following research question: How to design a computationally efficient data-driven weather forecasting system that learns the continuous latent representation while respecting the fundamental laws of atmospheric physics. In summary, our contributions are as follows:
  1. 1.
@@
    We propose continuous spatio-temporal continuous attention to effectively learn and interpolate the spatio-temporal information for weather forecasting.
  2. 2.
@@
    We introduce derivation as a pre-processing step to ensure better feature extraction for continuous spatio-temporal models.
  3. 3.
    We introduce physical constraints in our model via soft penalties in the form of a custom physics informed loss based on three fundamental physical laws of thermodynamics, kinetic enrgy and potential energy.
  4. 4.
    We perform extensive experiments on both WeatherBench and WeatherBench 2 to show the competitive performance of STC-ViT against state-of-the-art forecasting models.
@@ 2.1 Neural ODE
Neural ODEs are the continuous time models which learn the evolution of a system over time using Ordinary Differential Equations (ODE) (Chen et al., 2018). The key idea behind Neural ODE is to model the derivative of the hidden state using a neural network. Consider a hidden state $h(t)$ at time $t$ In a traditional neural network, the transformation from one layer to the next could be considered as moving from time $t$ to $t+1$. In Neural ODEs, instead of discrete steps, the change in $h(t)$ over time is defined by an ordinary differential equation parameterized by a neural network:
@@
ResNets vs Neural ODEs To see the connection between ResNets and Neural ODEs, consider a ResNet with layers updating the hidden state as:
@@
2.2 Physics Constrained Models
In weather and climate modeling, incorporating physical constraints ensures that the model adheres to the governing physical laws. These constraints can be added in two ways: hard constraints and soft constraints. Hard Constraints: For a given physical constraint $f(x)=0$ where $f(x)$ is the governing physical law, a hard constraint would mean that the machine learning model’s prediction $\hat{y}$ must always satisfy equation 4.
$f(\hat{y})=0$ (4)
This constraint can be embedded into model’s architecture as a constrained layer or optimizer.
Soft Constraints: Soft constraints adds a penalty term to the loss function that minimizes the violation of a physical law $f(\hat{y})$ as shown in equation 5
$min_{\theta}=L(y,\hat{y})+\alpha||f(\hat{y})||^{2}$ (5)
where $\alpha$ controls the weight of the penalty for violating the physical constraint $f(\hat{y})$ and $||f(\hat{y})||^{2}$ measures the degree of violation (e.g., deviation from mass conservation).
@@ 3 Related Work
WeatherBench by Rasp et al. (2020) provides a benchmark platform to evaluate data-driven systems for effective development of weather forecasting models. Current state-of-the-art data-driven forecasting models are mostly based on Graph Neural Networks (GNNs) and Transformers. Keisler (2022) implemented a message passing GNN based forecasting model which was further extended by GraphCast (Lam et al., 2023) which used multi-mesh GNN to achieve state-of-the-art results. FourCastNet (FCN) Kurth et al. (2023) used a combination of Fourier Neural Operator (FNO) and ViT and was reported to be 80,000 times faster than NWP models. Several more transformer based models emerged including Pangu-Weather (Bi et al., 2023), ClimaX (Nguyen et al., 2023a), FengWu (Chen et al., 2023a), FuXi (Chen et al., 2023b) and Stormer (Nguyen et al., 2023b) all showcased remarkable capabilities for short to medium range forecasting.
While being highly accurate and showcasing remarkable scaling capabilities, these models are discrete space-time models and do not account for the continuous dynamics of weather system. Recently, Verma et al. (2023) proposed ClimODE which used Neural ODE to incorporate a continuous-time process that models weather evolution and advection, enabling it to capture the dynamics of weather transport across the globe effectively. However it currently yields less precise forecast results compared to state-of-the-art models, offering significant potential for further enhancements. Further, Kochkov et al. (2023) proposed Neural GCM, which integrates a differentiable solver with neural networks resulting in physically consistent models.
@@ 4 Methodology
Problem Formulation. Consider a model $f$ receives weather data as input of the form $X^{N\times W\times H}$ at time $t$ and predicts the weather information at time $t+\triangle t$ as shown in equation 6 where $N$ is the number of weather variables such as temperature, wind speed, pressure, etc. and $H\times W$ refers to the spatial resolution of the variable.
@@
where $V(x,y,t)$ is pixel value of weather variable at time t and $V(x,y,t-1)$ is the pixel value of same variable at t-1.
@@ 4.1 Spatio-Temporal Continuous Vision Transformer
In the STC-ViT architecture, we enhance vision transformer-based weather forecasting by introducing Temporal Continuous Attention (TCA) and Spatial Attention (SA) mechanisms to capture dynamically evolving weather patterns. We build upon the variable tokenization and aggregation scheme proposed by (Nguyen et al., 2023a) followed by Continuous attention mechanism to enhance interpolated feature learning. The detailed architecture of STC-ViT is shown in figure 1.
Temporal Continuous Attention (TCA). We model the temporal dynamics with attention by incorporating derivatives (temporal changes) over time directly into the query, key, and value representations.
We formulate Query (Q) as the "current" time step’s information and is designed to seek relevant patterns or changes reflecting the model’s focus on upcoming changes equation 8.
$Q_{t}=\frac{dx_{t}}{dt}$ (8)
Key (K) models the "context" of prior time step states, which provide historical dynamics incorporating past time step derivatives equation 9.
$K_{t-1}=f(x_{t-1},\frac{dx}{dt}|_{t-1})$ (9)
Value (V) represents the updated information allowing the model to interpret them in the context of previous and current changes. This modification allows the model to learn the transitional changes from one time step to another which is important for capturing unprecedented changes in weather.
@@
To compute TCA and capture the temporal change for the same variable, we calculate the attention between $Q_{t_{0},i}$ and $K_{t-1,i}$ focusing on the temporal evolution. Assume we have two input samples of size $(N\times W\times H)$ denoted as $X_{t0}$ at time $t_{0}$ and $X_{t-1}$ at time $t_{-1}$ where $N$ is the number of weather variables such as temperature, wind speed, pressure, etc., $H\times W$ refers to the spatial resolution of the variable. The attention mechanism to capture the temporal continuity between time steps $t_{0}$ and $t_{-1}$ is given by:
$\frac{d}{dt}(Q_{t_{0},i}\cdot K_{t-1,i})=Q_{t_{0},i}\frac{d(K_{t-1,i})}{dt}+K_{t-1,i}\frac{d(Q_{t_{0},i})}{dt}$ (10)
$Y_{i}=Q_{t_{0},i}\frac{d(K_{t-1,i})}{dt}+K_{t-1,i}\frac{d(Q_{t_{0},i})}{dt}$ (11)
$Att(Q_{t_{0},i},K_{t-1,i})=Softmax(\frac{Y_{i}}{\sqrt{d_{k}}})$ (12)
$TCA_{t_{0},t-1}=\sum\limits^{N}_{k=1}Att(Q_{t_{0},i},K_{t-1,i})\cdot V_{t-1,i}$ (13)
@@
The resulting output represents the attention weighted sum of values for similar variables across input samples at time $t_{0}$ and $t_{1}$. This approach models the dynamics of each variable independently over time, allowing the model to capture temporal dependencies and changes effectively.
Spatial Attention (SA). For each variable i, we calculate the attention between spatial positions $(x,y)$ and $(x^{\prime},y^{\prime})$ given by equation 15:
$Attention(Q_{x,y,i},K_{x^{\prime},y^{\prime},i})=softmax(\frac{Q_{x,y,i}\cdot K_{x^{\prime},y^{\prime},i}}{\sqrt{d_{k}}})$ (14)
$SA_{t_{0},i}=\sum_{x^{\prime},y^{\prime}}Attention(Q_{x,y,i},K_{x^{\prime},y^{\prime},i})\cdot V_{x^{\prime},y^{\prime},i}$ (15)
where Q, K are matrices formed from all queries, keys, and values, respectively, $d_{k}$is the dimensionality of the keys and queries, and the division by $\sqrt{d_{k}}$ is a scaling factor to deal with gradient vanishing problems during training. Use the spatial attention weights to obtain the spatially-enhanced representation for each variable i by weighting the values:
@@
$STC=concat(TCA_{t_{0},i},SA_{t_{0},i})$ (16)
This output captures both the temporal evolution of each variable over time and the spatial interactions within each time step, enhancing the model’s ability to understand the continuous, spatiotemporal dynamics of the data.
@@
4.2 Latitude Weighted Physics Informed Loss Function
We use latitude weighted mean squared error to compute loss for the predicted variables.
$L_{lat-weight}=\frac{1}{V\times H\times W}\sum_{v=1}^{V}\sum_{i=1}^{H}\sum_{j=1}^{W}(L_{i})(\hat{X}_{t+\Delta t}^{n,i,j}-X_{t+\Delta t}^{n,i,j})$ (19)
where $L_{i}$ accounts for Latitude weights:
$L(i)=\frac{\cos(lat(i))}{\frac{1}{H}\sideset{}{{}_{i^{\prime}=1}^{H}}{\sum}\cos(lat(i^{\prime}))}$
Additionally, we account for three fundamental physical laws i.e. Potential Energy, Kinetic Energy and Thermodynamic Balance in our loss function. Kinetic energy in atmospheric science refers to the energy associated with the motion of air masses. Geopotential height is often used in meteorology to express the potential energy of an air parcel in the Earth’s gravitational field. It represents the energy per unit mass of an air parcel at height z above sea level and thermodynamic balance equation describes the evolution of temperature in an air parcel due to processes like heat addition and pressure changes. These fundamental principles, coupled with wind components and thermodynamic variables like temperature and geopotential height, form the core of atmospheric dynamics used in climate and weather models.
$L_{kinetic}=\left|\frac{1}{2}(u^{2}_{pred}+v^{2}_{pred})-\frac{1}{2}(u^{2}_{true}+v^{2}_{true})\right|$ (20)
where u is the eastward wind component (m/s) and v is the northward wind component (m/s).
$L_{potential}=|g\cdot z_{pred}-g\cdot z_{true}|$ (21)
where g is the acceleration due to gravity approximately $9.81m/s^{2}$ and z is the height or geopotential height (m).
$L_{thermo}=\left|\frac{dT}{dt}+u\frac{dT}{dx}+v\frac{dT}{dy}\right|$ (22)
where $\frac{dT}{dt}$ is the temporal change of temperature (model output), $u\frac{dT}{dx},v\frac{dT}{dy}$ represent the advection of temperature by wind in the x and y directions, respectively.
$Loss_{physics}=L_{lat-weight}+\alpha\times L_{kinetic}+\beta\times L_{potential}+\gamma\times L_{thermo}$ (23)
where $\alpha$, $\beta$ and $\gamma$ are the weight factors for physics based loss. The resulting loss aims to ensure that the predicted temperature and wind fields satisfy the physical laws of thermodynamics (like energy conservation) on their own. By setting up the equation to balance energy and temperature gradients, this loss constrains the predictions to maintain internal consistency with physical laws, regardless of the specific values in the ground truth.
Figure 1: Overall prediction pipeline of STC-ViT as the model receives spatio-temporal weather information and passed through pre-derivation and normalization steps and input to transformer encoder where TCA and SA learns the continuous weather features and outputs the prediction